OpenStudy (anonymous):

A linear equation in the form y = mx + b is written in slope-intercept form. What does the 'm' represent? A. x-intercept B. y-intercept C. z-intercept D. slope 2. A linear equation in the form y = mx + b is written in slope-intercept form. What does the 'b' represent? A. x-intercept B. y-intercept C. z-intercept D. slope 3. What is slope of equation y = 3x + 5? A. 3 B. -5 C. 5 D. -3 4. What is y-intercept of equation y = 3x + 5? A. 3 B. 5 C. -5 D. -3

OpenStudy (anonymous):

1. D 2. B I'm not to sure about 3 and for. Cause I think that 3 is -5 and 4 is 3

OpenStudy (anonymous):

5. Identify the slope and y-intercept of the line represented by the following linear equation: y = 2x + 3 A. slope = 1, y-intercept = 2 B. slope = 2, y-intercept = 3 C. slope = 3, y-intercept = 2 D. slope = 2, y-intercept = 1 6. What is first step in graphing using slope intercept form of an equation? A. Plug in 1 for y to find x. B. Place a point at origin then find second point using slope. C. Place the y-intercept on graph at point (0,b) 7. What is the second step in graphing using the slope-intercept form of an equation? A. From the y-intercept, count out the slope to place a second point on graph. B. Connect the two intercepts on the graph. C. Solve the equation for x.

OpenStudy (anonymous):

okay thank you any help is great

OpenStudy (anonymous):

6. B 7. A not to sure about 5. I think that is C

OpenStudy (anonymous):

Your welcome

OpenStudy (anonymous):

A. y = x - 5 B. y = 2/3x + 5 C. y = 2/3x - 5 D. y = x + 5

OpenStudy (anonymous):

the question is asking Which equation matches this graph?

OpenStudy (anonymous):

the picture is in the attachment